For 40 eruptions of the Old Faithful geyser in Yellowstone National Park, duration times (sec) were recorded along with the heights (ft) of the eruptions. The linear correlation coefficient is r = 0.0915 and the regression equation is y ̂ = 119 + 0.0331x, where x represents duration time. The mean of the 40 duration times is 245.0 sec and the mean of the 40 heights is 127.2 ft. What is the best predicted height of an eruption with a duration time of 120 min?
Answer #1

X = duration times (sec)

X = 120 min = 120*60 seconds

X = 7200 seconds

Model

y ̂ = 119 + 0.0331x

y ̂ = 119 + 0.0331*7200

y ̂ = 119 + 238.32

y ̂ = 357.32 ft.